### v1

# CAST: Contrastive Adaptation and Distillation for Semi-Supervised Instance Segmentation

Abstract
Instance segmentation demands costly per-pixel annotations and large models. We introduce CAST, a semi-supervised knowledge distillation (SSKD) framework that compresses pretrained vision foundation models (VFM) into compact experts using limited labeled and abundant unlabeled data. CAST unfolds in three stages: (1) domain adaptation of the VFM teacher(s) via self-training with contrastive pixel calibration, (2) distillation into a compact student via a unified multi-objective loss that couples standard supervision and pseudo-labels with our instance-aware pixel-wise contrastive term, and (3) fine-tuning on labeled data to remove residual pseudo-label bias. Central to CAST is an instance-aware pixel-wise contrastive loss that fuses mask and class scores to mine informative negatives and enforce clear inter-instance margins. By maintaining this contrastive signal across both adaptation and distillation, we align teacher and student embeddings and fully leverage unlabeled images. On Cityscapes and ADE20K, our 11$\approx$ 11$\times$ 11 × smaller student surpasses its adapted VFM teacher(s) by +3.4 AP (33.9 vs. 30.5) and +1.5 AP (16.7 vs. 15.2) and outperforms state-of-the-art semi-supervised approaches.
Appendix A Introduction
Pixel-level instance segmentation is notoriously expensive: annotating detailed masks can take hours per image, and training state-of-the-art detectors often requires hundreds of GPU hours, putting many applications out of reach [12, 15]. At the same time, recent advancements in VFMs [22, 20, 43, 19] have significantly elevated the capabilities of computer vision systems, enabling exceptional performance across diverse perception benchmarks [2].
Motivation. Despite remarkable achievements, foundation models still cannot serve specific downstream tasks sufficiently well due to two major issues: (1) the heavy computational overhead during deployment making these models impractical for environments with limited resources [37]; and (2) their inherently generic nature, which leads to suboptimal performance on tasks that demand domain specific expertise. The latter stems from foundation models being optimized to perform well across a wide variety of tasks, rather than being finely tuned for the nuanced requirements of specialized applications [5]. This challenge is particularly prominent in applications involving outdoor environments, such as autonomous driving, and indoor settings, such as robotic perception [13]. SSKD for instance segmentation seeks to compress large models into efficient student models by leveraging both limited labeled data and abundant unlabeled images. Current distillation methods either treat VFMs as fixed feature extractors with simple pseudo-labeling or focus on coarse semantic tasks, failing to exploit the rich structure of large unlabeled datasets to refine per-pixel predictions. Consequently, adjacent instances remain poorly separated and accuracy degrades sharply under scarce labels. We address these issues by adapting VFMs via self-training to enhance pseudo-label fidelity, and by injecting an instance-aware pixel-wise contrastive loss that leverages unlabeled data to enforce clear inter-instance margins, yielding sharper masks and superior performance in the low-label regime.
Status quo. Knowledge distillation has evolved from task-agnostic compression [16, 8] to adapting VFMs for downstream tasks. For classification and semantic segmentation, Vemulapalli et al. [30] distill a VFM matching its output on an unlabeled transfer set, and SAM-CLIP [31] fuses CLIP and SAM. However, neither method targets per-pixel instance masks nor exploits dense self-supervision from the unlabeled pool. Pure semi-supervised instance segmentation methods, such as [17, 3] train teachers from scratch, doubling GPU cost, and still produce blurred boundaries under scarce labels. To our knowledge, no prior work unifies VFM adaptation, unlabeled data-driven pixel-wise refinement, and extreme student compression for high fidelity instance segmentation.
Contributions. We summarize our main contributions as follows:
  1. 1.
    We introduce an instance-aware pixel-wise contrastive loss that fuses mask and class predictions to drive stronger inter-instance separation, and show how to sample negatives efficiently in a instance centric setting.
  2. 2.
    We propose CAST, a SSKD pipeline with three phases: (i) self-training a foundation teacher with contrastive calibration, (ii) distilling into a compact student using supervised, pseudo label, and pixel-wise contrastive objectives, and (iii) final supervised finetuning to remove bias thereby unifying supervised, semi-supervised, and self-supervised signals.
  3. 3.
    We conduct extensive experiments on Cityscapes and ADE20K, demonstrating that our 11$\approx$ 11$\times$ 11 × smaller student surpass self-trained teacher by +3.4 AP (33.9 vs. 30.5) and +2.1 AP (16.7 vs. 14.6), and outperforms state-of-the-art semi-supervised instance segmentation methods under the same data splits, with lower training cost.
Appendix B Related Work
Vision Foundation Models. Vision foundation models (VFMs)  [22, 20, 25, 41, 4] have revolutionized computer vision through large scale pretraining. In parallel, recent trends focus on combining VFMs to extend their capabilities [26, 43]. While these models excel in open-set recognition and transfer learning, their computational demands yet hinder edge deployment. Recent efforts merge VFMs via distillation: Wang et al [31] unify SAM and CLIP via multi-task learning, while Zhang etal [44] distill CLIP and DINOv2 into a compact model with moderate scale data distillation. We extend these paradigms by leveraging VFMs for instance segmentation, focusing on balancing robustness with computational efficiency.
Knowledge Distillation in Vision. Knowledge distillation (KD) has become a ubiquitous technique to transfer knowledge from teachers with high capacity to lightweight students for efficient deployment. Early methods distilled softened logits or intermediate features [16] in a task-agnostic way, while later feature-based approaches capture structured spatial cues (e.g., pixel-wise similarity, channel distributions) [23, 27]. Modern methods tackle VFMs’ scale and complexity: [28, 38] distills VFMs to impart zero-shot and multimodal capabilities, further multi teacher approaches [18, 39] combine complementary expertise. Vemulapalli et al. [30] adapt a VFM to the target task and then distill on a large unlabeled set for classification and semantic segmentation. Overall with these advances in vision KD We build our method on the insight that a strong teacher (or ensemble of teachers) can guide a lightweight instance segmentation model to high performance. our approach explicitly integrates semi-supervised learning and pixel-level contrastive losses for instance segmentation, to focus on bridging the gap between rich representation of VFMs and compact, efficient student networks.
Semi-Supervised Learning. Self-training (or pseudo-labeling) has become a foundational paradigm in semi-supervised learning (SSL), where a model leverages its own predictions with high confidence and iteratively refine itself [35]. This approach has proven effective across vision tasks, improving image classification performance [35] and boosting object detection accuracy when annotation budgets are tight [21]. To counteract error accumulation from noisy pseudo-labels [29] use exponential moving average of label predictions, or [6] employ curriculum labeling schemes that gradually incorporate harder examples. More recent work applies pseudo-labeling for large pretrained models through targeted finetuning and adaptive pseudo selection strategies [14]. While many SSL methods focus on classification or detection, several have extended this method to dense prediction tasks [10, 40].
We study self-training with self-supervised contrastive learning and task-specific adaptation. Global contrastive frameworks such as SimCLR [7], MoCo [9], and their detection extensions [34] established the value of large scale visual discrimination learning. Further per pixel contrastive approaches [32, 36, 46, 33, 1] have shown promise in retaining spatial sensitivity though they yet conflate pixels from different instances of the same class. We extend these advances by synergizing self-training and self-supervised contrastive learning, and introduce a novel instance aware negative sampling strategy designed specifically for the demands of instance segmentation.
Figure 1: CAST framework overview. Top: Three-stage pipeline: (1) adapt a pretrained VFM teacher to the target domain via self-training with contrastive pixel calibration; (2) transfer knowledge to a compact student using instance-aware pixel-wise contrastive sampling; (3) fine-tune the student on labeled data to correct residual pseudo-label bias. Bottom: Detailed view of stage (2): fused mask and class score maps produce anchor pixels, which are sampled across weak/strong views to form positive and negative pairs; an MLP projects features and computes the pixel-wise contrastive loss. Dashed arrows denote no gradient flow; red components are trainable, blue are frozen.
Appendix C Method
C.1 Overview
In semi-supervised settings, we are given a small labeled set and a substantially larger unlabeled pool:
$\mathcal{D}^{l}=\bigl{\{}(x_{i}^{l},y_{i}^{l})\bigr{\}}_{i=1}^{N_{l}}\quad\text{and}\quad\mathcal{D}^{u}=\bigl{\{}x_{i}^{u}\bigr{\}}_{i=1}^{N_{u}},\quad N_{u}\gg N_{l},$
where each $y_{i}^{l}$ consists of binary masks and class labels for every instance. Our goal is to distill knowledge from a large, pretrained VFM into a compact student $f_{\theta_{s}}$, matching or surpassing the teacher’s accuracy with far fewer labels and compute. We propose CAST, a three-stage SSKD pipeline that hinges on two core innovations: ❶ Contrastive Calibration. We fine-tune a large VFM teacher via self-training, but rather than simple pseudo-labels we inject a pixel-wise contrastive head to sharpen mask boundaries. ❷ Debiased, Instance-Aware Sampling. During both adaptation and distillation, we mine hard negatives via a joint mask-/class-probability embedding, focusing repulsion on informative inter-instance pairs tailored for instance segmentation. These two ideas are then realized in three concise stages (see Fig. 1):
  1. 1.
    Teacher Adaptation. Self-train the VFM with pseudo-labels and pixel-wise contrastive calibration to produce masks specialized to the target domain.
  2. 2.
    Knowledge Transfer. Freeze this calibrated teacher and distill into a lightweight student under a unified loss that harmonizes ground truth, pseudo-label, and contrastive terms, guided by our debiased sampling.
  3. 3.
    Student Refinement. Fine-tune the student on labeled data to remove residual pseudo-label bias.
Sec. C.2 formalizes our instance-aware pixel-wise contrastive loss, which is used in both Teacher Adaptation and Knowledge Transfer to enforce intra-instance cohesion and inter-instance separation; Sec. C.3 then details the three stages of the CAST pipeline.
C.2 Pixel-wise Contrastive Loss
Standard supervised and pseudo-label losses enforce correct mask predictions, ignoring pixel-level feature relationships which underutilize unlabeled data and amplify pseudo label noise. We therefore inject a self-supervised pixel-wise contrastive loss as an additional supervisory signal on both labeled and unlabeled images, sharpening feature discrimination and regularizing against noisy labels.
Let $z^{\rm weak},z^{\rm strong}\in\ R^{B\times N\times D}$ be $\ell_{2}$-normalized embeddings from two views of each image, where $B$ is the number of images in one mini batch, $N=h\times w$ the number of pixels, and $D$ the embedding dimension. For each pixel $p\in{1,2,...,N}$ and image index $b\in{1,...,B}$, the corresponding embedding vector is denoted as $z_{b,p}\in\mathbb{R}^{D}$. We construct the positive pair by sampling the weak and strong embeddings for each pixel. The positive similarity between the two views is
$s_{b,p}^{+}=\langle z^{\rm weak}_{b,p},\,z^{\rm strong}_{b,p}\rangle/T.$
Negatives are sampled by our instance-aware sampler (§C.2), producing indices $\{(b^{\prime},q_{r})\}_{r=1}^{R}$ and corresponding similarities $s^{-}_{b,p,r}$.
$s^{-}_{b,p,r}=\langle z^{\rm weak}_{b,p},\,z^{\rm strong}_{b^{\prime},q_{r}}\rangle/T,\quad r=1,\dots,R.$
The pixel-wise contrastive loss is then the standard NT-Xent over all anchors:
$\mathcal{L}_{\rm pxl}=-\frac{1}{B\,N}\sum_{b=1}^{B}\sum_{p=1}^{N}\log\frac{\exp(s^{+}_{b,p})}{\exp(s^{+}_{b,p})+\sum_{r=1}^{R}\exp(s^{-}_{b,p,r})}.$
Debiased Pixel-Level Negative Sampling.
To mine true inter-instance pairs without quadratic cost, we derive a per pixel sampling distribution by fusing mask and class probabilities. Let $M\in\mathbb{R}^{B\times K\times H\times W}$, and $L\in\mathbb{R}^{B\times K\times(C+1)},$ be the model’s mask and class logits respectively. We first resize $M$ to the feature resolution $(h\times w)$ and then normalize logits to probability distributions $P_{m}$ and $P_{c}$ via softmax along instance and class dimensions respectively.
For each pixel index $(b,p)$ to find the aggregated class vote, we compute Expected class distribution $F_{c}$. Further to avoid losing encoded instance ids over aggregation in expected class distribution we form a joint “pseudo probability” embedding by concatenation the mask distribution and class cues in a single vector which gives a richer embedding letting the contrastive head learn arbitrary interactions between mask and class. leading to pseudo probability map be $y[b,p]$.
$F_{c}[b,p,c]=\sum_{k=1}^{K}P_{m}[b,k,p]\;P_{c}[b,k,c],\quad y[b,p]=\begin{bmatrix}P_{m}[b,1:K,p]\\[6.0pt] F_{c}[b,p,1:C+1]\end{bmatrix}\;\in\;\ R^{\,K+(C+1)}.$
We score any two pixels $(b,p)\neq(b^{\prime},q)$ by $\tilde{y}$ being $\ell_{2}$-normalized vector of pseudo probability map.
[…]
We draw $R$ negatives $\{q_{r}\}$ for each anchor $(b,p)$ by sampling proportional to $s^{\rm deb}$, and then plug these into the NT-Xent denominator of $\mathcal{L}_{\rm pxl}$.
Theoretical Insight. To give a formal rationale for augmenting our pixel-wise contrastive loss, we show that even under a mild negative sampling guarantee, each gradient step on our contrastive term provably increases the expected inter-instance margin.
Assumption C.1 (Negative Sampling Guarantee).
When sampling a negative under our instance aware scheme, the probability it originates from a different instance is at least $p>0.5$, where $p$ can be estimated empirically (see Sec. D.3).
Proposition C.1 (Expected Margin Growth).
Under Assumption C.1, one gradient update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin $\Delta_{\rm emp}$ by
$\varepsilon=\Theta(p\,\lambda_{\rm pxl})>0.$
This expectation holds even when pseudo-labels are imperfect, provided negatives are sampled using our instance aware strategy.
In practice, raising $\lambda_{\rm pxl}$ enhances margin growth but also increases training cost. If $\lambda_{\rm pxl}$ is too large, it can overemphasize inter-instance separation at the expense of intra-instance cohesion. We validate this effect in Sec. D.3 and provide a proof sketch in Appendix C.
C.3 CAST Framework
We cast teacher adaptation, student distillation and student refinement as special cases of the same objective with three terms. Let
$\mathcal{J}(\theta;\,\mathcal{D}^{l},\mathcal{D}^{u};\,\lambda_{\rm semi},\lambda_{\rm pxl})=\underbrace{\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\ell\bigl{(}f_{\theta}(x_{i}^{l}),y_{i}^{l}\bigr{)}}_{\mathcal{L}_{\rm sup}}+\lambda_{\rm semi}\underbrace{\frac{1}{N_{u}}\sum_{j=1}^{N_{u}}\ell\bigl{(}f_{\theta}(x_{j}^{u}),\hat{y}_{j}^{u}\bigr{)}}_{\mathcal{L}_{\rm semi}}+\lambda_{\rm pxl}\,\mathcal{L}_{\rm pxl}\bigl{(}\theta;\,\mathcal{D}^{l}\cup\mathcal{D}^{u}\bigr{)},$
where $\mathcal{D}^{u}=\varnothing$ makes the middle term zero.
Teacher adaptation. Starting from pretrained weights $\theta_{T}^{0}$, we first fine-tune on the labeled set $\mathcal{D}^{l}$:
[…]
We then generate pseudo-labels $\hat{y}_{j}^{u}=f_{\theta_{T}^{\prime}}(x_{j}^{u})$, reset to $\theta_{T}^{0}$and fine-tune on $\mathcal{D}^{l}\cup\{\!(x_{j}^{u},\hat{y}_{j}^{u})\!\}$:
[…]
This two-step contrastive calibration yields a specialized teacher whose pseudo-labels are both accurate and spatially consistent for the target domain.
Knowledge transfer. With calibrated teacher $\theta_{T}^{\prime\prime}$ frozen, student $\theta_{s}$ is trained via the unified objective:
$\theta_{s}^{*}=\arg\min_{\theta_{s}}\;\mathcal{J}\bigl{(}\theta_{s};\,\mathcal{D}^{l},\mathcal{D}^{u};\,\lambda_{\rm semi},\lambda_{\rm pxl}\bigr{)}.$ (1)
Here, $\mathcal{L}_{\rm sup}$ enforces ground truth supervision on $\mathcal{D}^{l}$, $\mathcal{L}_{\rm semi}$ distills pseudo-labels from $\mathcal{D}^{u}$, and $\mathcal{L}_{\rm pxl}$ imposes our pixel-wise contrastive regularizer across both sets. The coefficients $\lambda_{\rm semi}$ and $\lambda_{\rm pxl}$ balance signals, guiding the student to approach teacher’s accuracy with far fewer parameters.
Student Refinement. Although joint distillation yields a strong initialization, residual pseudo-label noise and contrastive pretext tasks can introduce bias. As a final step, we fine-tune the student on labeled data alone:
[…]
This pass removes pseudo-label drift and sharpens decision boundaries for in-domain data.
Appendix D Experiments
D.1 Experimental Protocol
Datasets. We evaluate CAST on two standard instance segmentation benchmarks: Cityscapes [12] contains 2,975 training, 500 validation images of urban street scenes, annotated with 19 semantic categories (8 “thing” classes and 11 “stuff” classes). ADE20K [47] comprises 20,210 training and 2,000 validation images spanning diverse indoor and outdoor environments, annotated with 150 semantic categories (100 “thing” and 50 “stuff” classes).
For our semi-supervised experiments, we randomly sample $10\%$ of the Cityscapes training set as labeled and use the remaining images as unlabeled. For ADE20K, to reduce computational cost we use stratified sampling to select $20\%$ of the training pool while preserving the original per category distribution, yielding 3,537 images. From this subset, 1,000 images (10 per instance class) serve as labeled data and the remaining 2,537 as unlabeled. Detailed statistics are provided in Appendix A, and we will release the exact split files alongside our code.
Teacher and Student Architectures. Our teacher is a fused ensemble of Grounding-DINO-Large [20] and SAM2-L [25]. Since the official Grounding-DINO is closed-source, we use its open-source counterpart mm-Grounding-DINO [45]. For the student, we pair a DINOv2-S encoder [22] with a DPT-S decoder head [24], followed by a lightweight transformer decoder module in the spirit of Mask2Former [11]. Our choice of the DINOv2+DPT backbone is motivated by the recent successes of “Depth AnythingV2” in monocular depth estimation [41] and UniMatchV2 in semantic segmentation [42], and aims to facilitate future multimodal fusion work. We evaluate the impact of different student designs in Sec. 4, and defer the complete optimizer, learning rate schedules, and other hyperparameters to Appendix B.
Implementation Details. All experiments were conducted on Ubuntu 22.04 with Python 3.10 and PyTorch 2.6.0 (CUDA 12.6). Teacher adaptation runs were executed on 2× NVIDIA A100 GPUs, while student training and fine-tuning runs used 2× NVIDIA GeForce RTX 4090 GPUs. For instance, a single fine-tuning run of the teacher (Grounding-DINO) on the supervised Cityscapes split required $\approx 3.5$ GPU hours; a single student training run for this dataset took $\approx 17$ GPU hours without the pixel-wise contrastive branch and $\approx 22$ GPU hours with it. All reported results in the following tables are averages over three independent runs with different random seeds.
D.2 Main Results
We extensively evaluate a range of knowledge distillation (KD) strategies, ranging from purely supervised to state-of-the-art semi-supervised baselines, and benchmark them against our CAST pipeline. Table 1 reports maskAP and AP50 on Cityscapes and ADE20K. In the teacher adaptation stage (568 M parameters), adding our pixel-level contrastive loss boosts Cityscapes maskAP from 29.8 to 30.5 (+0.7) and AP50 from 54.9 to 56.6 (+1.7); on ADE20K, maskAP rises from 14.6 to 15.2 (+0.6) and AP50 from 23.6 to 24.5 (+0.9). These improvements confirm that pixel-wise supervision sharply improves feature discrimination and reduces pseudo-label noise. Figure 2 plots key pipeline efficiency metric on a logarithmic scale for both teacher and student models.
Figure 2: Comparison of Teacher and Student Model efficiency Metrics (Log Scale).
In the student distillation stage (52 M parameters, just 9 % of the teacher), our semi-supervised model with pixel-level loss achieves 32.2 maskAP and 56.5 AP50 on Cityscapes, outperforming the Guided dist. baseline by +1.4 maskAP. After fine-tuning, the student reaches 33.9 maskAP (+3.4 over the best teacher) and 58.7 AP50. On ADE20K, it records 16.1 maskAP and 27.4 AP50 semi-supervised, and 16.7 maskAP (+1.5) with 28.0 AP50 after fine-tuning, underscoring CAST’s robustness across domains.
Table 1: Instance segmentation results on Cityscapes and ADE20K. We report teacher-adaptation (568 M params) and student distillation (52 M params) under each method.
Method Data Regime Cityscapes ADE20K
maskAP AP50 maskAP AP50
Teacher Adaptation
[…]
Supervised Labeled only $28.7$ $53.4$ $14.2$ $23.5$
Self-training [35] Labeled + Unlabeled $29.7$ $54.9$ $14.6$ $23.6$
+ EMA-Augmented [21] Labeled + Unlabeled $29.8$ $54.9$ $14.8$ $23.7$
+ pixel-level loss (ours) Labeled + Unlabeled 30.5 56.6 15.2 24.5
Student Distillation
[…]
PAIS [17] Labeled + Unlabeled $22.9$ $44.9$ $10.3$ $18.3$
Guided dist. [3] Labeled + Unlabeled $30.8$ $52.9$ $14.2$ $23.8$
+ unlabeled KD [30] Unlabeled only $24.4$ $45.6$ $5.1$ $9.3$
+ labeled+unlabeled KD (ours) Labeled + Unlabeled $30.7$ $54.9$ $14.4$ $25.2$
+ pixel-level loss (ours) Labeled + Unlabeled 32.2 56.5 16.1 27.4
Fine-tuned (best) Labeled only 33.9 58.7 16.7 28.0
D.3 Empirical Validation
We validate Proposition C.1 by monitoring the false negative rate ($\mathrm{FNR}$), the fraction of sampled negatives that actually belong to the same instance, and the empirical margin
$\Delta_{\rm emp}=\mathrm{NegMean}-\mathrm{PosMean}.$
Defining $p=1-\mathrm{FNR}$ as the success probability of sampling a true negative, Figure 3 shows: the empirical margin every 10 k iterations for $\lambda_{\rm pxl}\in\{0.01,0.05,0.1,0.2\}$ (left), the raw contrastive loss for $\lambda_{\rm pxl}=0.1$ (center), and the false negative rate for $\lambda_{\rm pxl}=0.1$ (right, dashed at $p=0.5$). Throughout training we observe $p>0.9$ and a linear increase of $\Delta_{\rm emp}$ with $\lambda_{\rm pxl}$, in agreement with Proposition C.1.
(a) Instance Separation
(b) $\mathrm{FNR}$
(c) Contrastive Loss
Figure 3: (Left) Empirical margin (NegMean–PosMean) every 10k iterations for various $\lambda_{\rm pxl}$. (Center) False negative rate ($\mathrm{FNR}$) for $\lambda_{\rm pxl}=0.1$, dashed at $p=0.5$. (Right) Contrastive loss for $\lambda_{\rm pxl}=0.1$.
D.4 Ablation Studies
We perform a series of ablation experiments on the Cityscapes dataset to isolate the contributions of different components in the CAST pipeline and the pixel-level contrastive loss.
Impact of Loss Components. In the knowledge distillation stage, the overall loss is a weighted sum of three components: the supervised loss ($\mathcal{L}_{\text{sup}}$), the semi-supervised pseudo-label loss ($\mathcal{L}_{\text{semi}}$), and the pixel-level self-supervised contrastive loss ($\mathcal{L}_{\text{pxl}}$). Table 2 presents an ablation study assessing the contribution of each loss term. The results indicate that adding the pseudo-label loss substantially improves performance over the supervised baseline, and further inclusion of the self-supervised loss leads to the highest overall performance.
Table 2: Loss Ablation: Impact of supervised ($\mathcal{L}_{\rm sup}$), pseudo-label ($\mathcal{L}_{\rm semi}$), and pixel-level ($\mathcal{L}_{\rm pxl}$) terms on teacher vs. student maskAP.
Method $\mathcal{L}_{\rm sup}$ $\mathcal{L}_{\rm semi}$ $\mathcal{L}_{\rm pxl}$ maskAP (%)
Teacher Student
(a) Supervised Only 28.7 21.1
(b) + Pseudo-Label 29.7 30.7
(c) + pixel-level loss 29.6 27.5
(d) + (b) + (c) 30.5 32.2
Ablation of Negative Sampling via Various Probability Maps.
To validate our negative sampling strategy in the pixel-level contrastive loss, Table 3 compares four sampling methods: Uniform: negatives sampled uniformly across the image; Mask-Only: The probability map is derived solely from mask predictions, with class probabilities assumed to be uniform. Class-Only: The map is generated only from class predictions, assuming a uniform spatial distribution for the mask. Fusion: Combining both mask and class predictions. The fusion strategy achieves the best results, with 32.2 maskAP and 56.5 AP50.
(a) Negative Sampling Strategies
Method maskAP (%) AP50 (%)
Uniform $29.4$ $50.2$
Mask-Only $30.6$ $55.0$
Class-Only $31.1$ $55.3$
Fusion 32.2 56.5
(b) Schematic of Sampling Distributions
Table 3: Ablation of Negative Sampling Strategies on Cityscapes. (3) Quantitative results for uniform, mask-only, class-only, and fusion samplers (maskAP and AP50). (3) Schematic sketch of the corresponding pixel-level sampling probability distributions.
Hyperparameter Sensitivity. We evaluate CAST’s sensitivity to three key hyperparameters on Cityscapes: contrastive weight $\lambda_{\rm pxl}$, negatives per anchor $K$, and temperature $T$, by measuring both teacher and student mask AP (%) and AP50 (%). Table 4 reports the full sweep. We find that $\lambda_{\rm pxl}=0.2$ and $T=0.2$ consistently maximize performance. For the number of negatives, $K=256$ offers the best trade-off: although $K=512$ yields a slight increase in teacher maskAP (30.9 vs. 30.5) and AP50 (57.1 vs. 56.6), and comparable student metrics, the marginal gains saturate relative to the increased sampling cost. Therefore, we adopt $K=256$ throughout.
Table 4: Hyperparameter Ablation on Cityscapes.
Model Metric Contrastive Loss Weight($\lambda_{\rm pxl}$) Negative Sanples per Anchor($K$) Temperature($T$)
0 0.01 0.1 0.2 0.5 128 256 512 0.1 0.2 0.4
Teacher AP $29.7$ $29.9$ $30.2$ $30.5$ $30.1$ $30.4$ $30.5$ $30.9$ $30.1$ $30.5$ $29.8$
AP50 $55.3$ $55.7$ $56.1$ $56.6$ $56.1$ $56.3$ $56.6$ $57.1$ $55.9$ $56.6$ $55.3$
Student AP $30.7$ $30.8$ $32.1$ $32.2$ $30.9$ $29.8$ $32.2$ $32.1$ $31.9$ $32.2$ $31.7$
AP50 $54.9$ $55.2$ $56.2$ $56.5$ $55.7$ $55.3$ $56.5$ $56.6$ $56.0$ $56.5$ $55.8$
Student Architecture Variants.
We evaluate two design axes for the student model under CAST distillation protocol: (i) the encoder backbone (with a fixed DPT decoder), and (ii) the decoder head (with a fixed DINOv2-S encoder). Table 5 reports accuracy along with parameter counts, on the Cityscapes validation set. The combination of DINOv2-S encoder and DPT head achieves the best accuracy with a compact footprint.
Table 5: Architecture Ablations on Cityscapes. (5(a)) Encoder backbone (fixed DPT decoder). (5(b)) Decoder head (fixed DINOv2-S encoder).
(a) Encoder Backbone
Encoder maskAP AP50 Params (M)
ResNet50 25.5 49.3 24
SAM2-S 22.1 39.2 35
DINOv2-S 30.7 54.9 22
(b) Decoder Head
Decoder maskAP AP50 Params (M)
FPN 28.9 52.4 18
DPT 30.7 54.9 22
Figure 4: Qualitative results on Cityscapes. Guided dist. [3] (top) vs. CAST (bottom).
Appendix E Conclusions
We have introduced CAST, a rigorously designed SSKD pipeline that fuses self-training, instance-aware pixel-wise contrastive learning, and final supervised finetuning to compress large VFMs into compact student experts with comparable performance. Empirically, our 11$\approx$ 11$\times$ 11 × smaller student exceeds its adapted teacher by +3.4 maskAP in Cityscapes and +1.5 maskAP in ADE20K, while cutting compute and parameter counts demonstrating that dense contrastive supervision can unlock substantial gains in low-label regimes. Our theoretical analysis further guarantees that our negative sampling scheme provably increases inter-instance margins under mild assumptions. Looking forward, streamlining CAST into a single unified objective, extending its evaluation to diverse domains, and integrating uncertainty quantification will be critical steps toward safe, equitable, and broadly deployable segmentation solutions.
Figure 5: Performance–complexity radar chart (normalized).
Broader Impacts. This paper focuses on developing high fidelity instance segmentation models under tight annotation and compute budgets, running efficiently on workstation GPUs and making it practical for domains such as autonomous driving, mobile and industrial robotics, augmented reality and agricultural or infrastructure monitoring. These safety critical applications are especially vulnerable to biases inherited from generalized VFMs (e.g., underrepresented object classes or ambiguous category definitions, which can lead to degraded performance or unfair outcomes. By distilling task-specific expert models, CAST helps reduce these generic failure modes, but careful domain specific validation, bias auditing, and uncertainty quantification remain essential for deployment. Limitations. While our CAST framework achieves significant performance improvements, the training pipeline comprising multiple stages introduces additional complexity. Future work could focus on streamlining these stages and further reducing the training cost.
Figure 6: Qualitative results on ADE20k.
Figure 7: Qualitative bias reduction in stage-wise distillation. Top row: pseudo-labels generated by the pretrained teacher. Bottom row: student predictions after distillation and refinement, demonstrating reduced pseudo-label bias and sharper instance boundaries.
Supplementary Material
This document provides additional details to support the main paper, including dataset statistics, full hyperparameter settings, formal proof, extended training protocols, and additional ablation studies.
Appendix F Dataset Splits
Table 6 summarizes the datasets used in our experiments. We use a 10% labelled split of Cityscapes’ 2 975 training images (298 labeled / 2 677 unlabeled) and a stratified 20% split of ADE20K’s 20 210 training images (1 000 labeled / 2 537 unlabeled). Standard validation sets are retained (500 images for Cityscapes, 2 000 for ADE20K). Exact image-ID lists will be released with our code.
Table 6: Semi-supervised splits used in our experiments.
Dataset # Classes Labeled / Unlabeled Validation
Cityscapes 8 298 / 2 677 500
ADE20K 100 1 000 / 2 537 2 000
Appendix G Hyperparameters
Key teacher and student hyperparameters are summarized in Table 7.
Table 7: Hyperparameter Settings
Parameter Teacher Student
Learning rate $5.0\times 10^{-5}$ Encoder: $5.0\times 10^{-6}$; Decoder: $5.0\times 10^{-5}$
Scheduler Multi-step (milestones at 0.9, 0.95) PolyLR (power 0.9)
Batch size 4 8
Weight decay 0.01 0.05
Contrastive loss weight 0.2 0.2
Pseudo-label threshold 0.3 0.3
Dropout rate 0.1
Gradient clipping $\ell_{2}$ norm 0.1
Optimizer AdamW ($\beta 1$=0.9, $\beta 2$=0.999)
Augmentations Weak: flip, resize; Strong: random resized crop, jitter, grayscale, blur,
Loss weights (mask / class) 5 / 2
Appendix H Proof Sketch of Proposition 3.1
Proof Sketch.
Let $z_{a}$, $z^{+}$ and $\{z^{-}_{r}\}_{r=1}^{R}$ be the unit norm embeddings of an anchor pixel, its positive, and $R$ negatives. Define
$s^{+}=\langle z_{a},\,z^{+}\rangle,\qquad s^{-}_{r}=\langle z_{a},\,z^{-}_{r}\rangle,$
and the pixel-wise contrastive loss
$\ell(z_{a})=-\log\frac{\exp(s^{+})}{\exp(s^{+})+\sum_{r=1}^{R}\exp(s^{-}_{r})}.$
Let
$Z=\exp(s^{+})+\sum_{r=1}^{R}\exp(s^{-}_{r}),\qquad\alpha_{r}=\frac{\exp(s^{-}_{r})}{Z}.$
A straightforward gradient computation gives
$\nabla_{z_{a}}\ell=\sum_{r=1}^{R}\alpha_{r}\,(z^{-}_{r}-z^{+}).$
Applying one gradient descent step with step size $\lambda_{\rm pxl}$:
$z_{a}^{\prime}=z_{a}-\lambda_{\rm pxl}\,\nabla_{z_{a}}\ell\;=\;z_{a}+\lambda_{\rm pxl}\sum_{r=1}^{R}\alpha_{r}\,(z^{+}-z^{-}_{r}).$
For a randomly chosen negative $z^{-}$,
$\Delta s^{+}$ $=\langle z_{a}^{\prime}-z_{a},\,z^{+}\rangle=\lambda_{\rm pxl}\sum_{r=1}^{R}\alpha_{r}\bigl{(}1-\langle z^{-}_{r},\,z^{+}\rangle\bigr{)},$
$\Delta s^{-}$ $=\langle z_{a}^{\prime}-z_{a},\,z^{-}\rangle=\lambda_{\rm pxl}\sum_{r=1}^{R}\alpha_{r}\bigl{(}\langle z^{+},\,z^{-}\rangle-\langle z^{-}_{r},\,z^{-}\rangle\bigr{)}.$
By Assumption 3.1, each negative embedding $z^{-}_{r}$ is inter-instance with probability $p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 0$, and intra-instance with probability $1-p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 1$. Hence
$\mathbb{E}\bigl{[}1-\langle z^{-}_{r},\,z^{+}\rangle\bigr{]}=p\cdot 1+(1-p)\cdot 0=p,$
and since $\sum_{r=1}^{R}\alpha_{r}=1$, it follows that
$\mathbb{E}[\Delta s^{+}]=\lambda_{\rm pxl}\sum_{r=1}^{R}\alpha_{r}\,\mathbb{E}\bigl{[}1-\langle z^{-}_{r},\,z^{+}\rangle\bigr{]}=p\,\lambda_{\rm pxl}.$
Meanwhile, every term in $\Delta s^{-}$ involves an inter-instance inner product, either $\langle z^{+},z^{-}\rangle$ or $\langle z^{-}_{r},z^{-}\rangle$ each of which vanishes in expectation, so $\mathbb{E}[\Delta s^{-}]\approx 0$. Therefore
$\mathbb{E}[\Delta s^{+}-\Delta s^{-}]=p\,\lambda_{\rm pxl}-0=\Theta\!\bigl{(}p\,\lambda_{\rm pxl}\bigr{)}=\varepsilon>0,$
i.e. one update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin by $\varepsilon$. ∎
Remark H.1 (Why $\langle z^{+},z^{-}\rangle\approx 0$ holds).
Under the InfoNCE objective (§3.2), the normalized weights for negative pairs, $\alpha_{r}=\frac{e^{s^{-}_{r}}}{e^{s^{+}}+\sum_{r}e^{s^{-}_{r}}},$ vanish at convergence, i.e. $\alpha_{r}\approx 0$. Moreover, in high dimensional embeddings, random unit vectors have inner products concentrating near zero, and contrastive training further pushes these negative similarities into a tight, small magnitude distribution [7]. Thus it is reasonable to approximate $\langle z^{+},z^{-}\rangle\approx 0$ up to $O(1/\sqrt{D})$ fluctuations.
Appendix I More Training Details
All teacher models are fine-tuned using 1k iterations on labeled set, followed by 5k iterations in a self-training stage with pseudo-labels. For student models, training on the Cityscapes dataset spans 90k iterations (consistent with prior work [11]), while the mini-ADE20k dataset is trained for 80k iterations. Finally, both datasets undergo an additional supervised fine-tuning phase for 2k iterations.
Appendix J Additional Ablations
J.1 Loss Variant: InfoNCE vs. Margin Hinge
Replacing our asymmetric InfoNCE (§3.2) with an margin-based hinge loss (margin = 0.2) yields identical maskAP (32.2%) and +0.6 AP50, at the cost of $1.6$× longer training. This evaluates whether enforcing a fixed positive–negative margin can match or improve upon the performance of InfoNCE.
Table 8: Loss Variant Ablation. Default InfoNCE vs. margin-based hinge (m = 0.2).
Loss Variant maskAP (%) AP50 (%)
Asymmetric InfoNCE (§3.2) 32.2 56.5
Margin hinge (m = 0.1) 32.2 57.1
J.2 Ablation: Debias Score Formulation
[…]
Table 9: Debias Score Formulation Ablation. Original vs. squared vs. square-root debias scores.
Score Variant maskAP (%) AP50 (%)
Original 32.2 56.5
Squared 32.0 56.3
Square-root 31.9 56.2
J.3 Ablation: Negative Sampling Scope
We evaluate two negative sampling scopes: (i) sampling only within the current mini batch vs. (ii) sampling from a small memory bank of past pixel embeddings (size 10k).
Table 10: Sampling Scope Ablation. Mini batch only vs. memory bank negatives.
Scope maskAP (%) AP50 (%)
Mini-batch only 32.2 56.5
Memory bank (10k embeddings) 32.7 57.3
Sampling from a memory bank of 10 k embeddings yields a modest performance gain (+0.5 maskAP, +0.8 AP50) compared to in-batch sampling. However, this approach incurs approximately $2.2$× longer training time due to the overhead of maintaining and querying the memory bank.

### v3

@@
 15.2) and outperforms state-of-the-art semi-supervised approaches.
-Appendix A Introduction
+Appendix 1 Introduction
 Pixel-level instance segmentation is notoriously expensive: annotating detailed masks can take hours per image, and training state-of-the-art detectors often requires hundreds of GPU hours, putting many applications out of reach [12, 15].
@@
 14.6), and outperforms state-of-the-art semi-supervised instance segmentation methods under the same data splits, with lower training cost.
-Appendix B Related Work
+Appendix 2 Related Work
 Vision Foundation Models.
@@
 Dashed arrows denote no gradient flow; red components are trainable, blue are frozen.
-Appendix C Method
-C.1 Overview
+Appendix 3 Method
+3.1 Overview
 In semi-supervised settings, we are given a small labeled set and a substantially larger unlabeled pool:
@@
 Sec.
-C.2 formalizes our instance-aware pixel-wise contrastive loss, which is used in both Teacher Adaptation and Knowledge Transfer to enforce intra-instance cohesion and inter-instance separation; Sec.
-C.3 then details the three stages of the CAST pipeline.
-C.2 Pixel-wise Contrastive Loss
+3.2 formalizes our instance-aware pixel-wise contrastive loss, which is used in both Teacher Adaptation and Knowledge Transfer to enforce intra-instance cohesion and inter-instance separation; Sec.
+3.3 then details the three stages of the CAST pipeline.
+3.2 Pixel-wise Contrastive Loss
 Standard supervised and pseudo-label losses enforce correct mask predictions, ignoring pixel-level feature relationships which underutilize unlabeled data and amplify pseudo label noise.
@@
 $s_{b,p}^{+}=\langle z^{\rm weak}_{b,p},\,z^{\rm strong}_{b,p}\rangle/T.$
-Negatives are sampled by our instance-aware sampler (§C.2), producing indices $\{(b^{\prime},q_{r})\}_{r=1}^{R}$ and corresponding similarities $s^{-}_{b,p,r}$.
+Negatives are sampled by our instance-aware sampler (§3.2), producing indices $\{(b^{\prime},q_{r})\}_{r=1}^{R}$ and corresponding similarities $s^{-}_{b,p,r}$.
 $s^{-}_{b,p,r}=\langle z^{\rm weak}_{b,p},\,z^{\rm strong}_{b^{\prime},q_{r}}\rangle/T,\quad r=1,\dots,R.$
@@
 To give a formal rationale for augmenting our pixel-wise contrastive loss, we show that even under a mild negative sampling guarantee, each gradient step on our contrastive term provably increases the expected inter-instance margin.
-Assumption C.1 (Negative Sampling Guarantee).
-When sampling a negative under our instance aware scheme, the probability it originates from a different instance is at least $p>0.5$, where $p$ can be estimated empirically (see Sec. D.3).
-Proposition C.1 (Expected Margin Growth).
-Under Assumption C.1, one gradient update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin $\Delta_{\rm emp}$ by
+Assumption 3.1 (Negative Sampling Guarantee).
+When sampling a negative under our instance aware scheme, the probability it originates from a different instance is at least $p>0.5$, where $p$ can be estimated empirically (see Sec. 4.3).
+Proposition 3.1 (Expected Margin Growth).
+Under Assumption 3.1, one gradient update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin $\Delta_{\rm emp}$ by
 $\varepsilon=\Theta(p\,\lambda_{\rm pxl})>0.$
@@
 We validate this effect in Sec.
-D.3 and provide a proof sketch in Appendix C.
-C.3 CAST Framework
+4.3 and provide a proof sketch in Appendix C.
+3.3 CAST Framework
 We cast teacher adaptation, student distillation and student refinement as special cases of the same objective with three terms. Let
@@
 This pass removes pseudo-label drift and sharpens decision boundaries for in-domain data.
-Appendix D Experiments
-D.1 Experimental Protocol
+Appendix 4 Experiments
+4.1 Experimental Protocol
 Datasets.
@@
 All reported results in the following tables are averages over three independent runs with different random seeds.
-D.2 Main Results
+4.2 Main Results
 We extensively evaluate a range of knowledge distillation (KD) strategies, ranging from purely supervised to state-of-the-art semi-supervised baselines, and benchmark them against our CAST pipeline.
@@
 Fine-tuned (best) Labeled only 33.9 58.7 16.7 28.0
-D.3 Empirical Validation
-We validate Proposition C.1 by monitoring the false negative rate ($\mathrm{FNR}$), the fraction of sampled negatives that actually belong to the same instance, and the empirical margin
+4.3 Empirical Validation
+We validate Proposition 3.1 by monitoring the false negative rate ($\mathrm{FNR}$), the fraction of sampled negatives that actually belong to the same instance, and the empirical margin
 $\Delta_{\rm emp}=\mathrm{NegMean}-\mathrm{PosMean}.$
 Defining $p=1-\mathrm{FNR}$ as the success probability of sampling a true negative, Figure 3 shows: the empirical margin every 10 k iterations for $\lambda_{\rm pxl}\in\{0.01,0.05,0.1,0.2\}$ (left), the raw contrastive loss for $\lambda_{\rm pxl}=0.1$ (center), and the false negative rate for $\lambda_{\rm pxl}=0.1$ (right, dashed at $p=0.5$).
-Throughout training we observe $p>0.9$ and a linear increase of $\Delta_{\rm emp}$ with $\lambda_{\rm pxl}$, in agreement with Proposition C.1.
+Throughout training we observe $p>0.9$ and a linear increase of $\Delta_{\rm emp}$ with $\lambda_{\rm pxl}$, in agreement with Proposition 3.1.
 (a) Instance Separation
@@
 (Right) Contrastive loss for $\lambda_{\rm pxl}=0.1$.
-D.4 Ablation Studies
+4.4 Ablation Studies
 We perform a series of ablation experiments on the Cityscapes dataset to isolate the contributions of different components in the CAST pipeline and the pixel-level contrastive loss.
@@
 Figure 4: Qualitative results on Cityscapes. Guided dist. [3] (top) vs. CAST (bottom).
-Appendix E Conclusions
+Appendix 5 Conclusions
 We have introduced CAST, a rigorously designed SSKD pipeline that fuses self-training, instance-aware pixel-wise contrastive learning, and final supervised finetuning to compress large VFMs into compact student experts with comparable performance.
@@
 This document provides additional details to support the main paper, including dataset statistics, full hyperparameter settings, formal proof, extended training protocols, and additional ablation studies.
-Appendix F Dataset Splits
+Appendix 6 Dataset Splits
 Table 6 summarizes the datasets used in our experiments.
@@
 ADE20K 100 1 000 / 2 537 2 000
-Appendix G Hyperparameters
+Appendix 7 Hyperparameters
 Key teacher and student hyperparameters are summarized in Table 7.
@@
 Loss weights (mask / class) 5 / 2
-Appendix H Proof Sketch of Proposition 3.1
+Appendix 8 Proof Sketch of Proposition 3.1
 Proof Sketch.
@@
 i.e. one update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin by $\varepsilon$. ∎
-Remark H.1 (Why $\langle z^{+},z^{-}\rangle\approx 0$ holds).
+Remark 8.1 (Why $\langle z^{+},z^{-}\rangle\approx 0$ holds).
 Under the InfoNCE objective (§3.2), the normalized weights for negative pairs, $\alpha_{r}=\frac{e^{s^{-}_{r}}}{e^{s^{+}}+\sum_{r}e^{s^{-}_{r}}},$ vanish at convergence, i.e.
@@
 Thus it is reasonable to approximate $\langle z^{+},z^{-}\rangle\approx 0$ up to $O(1/\sqrt{D})$ fluctuations.
-Appendix I More Training Details
+Appendix 9 More Training Details
 All teacher models are fine-tuned using 1k iterations on labeled set, followed by 5k iterations in a self-training stage with pseudo-labels.
@@
 Finally, both datasets undergo an additional supervised fine-tuning phase for 2k iterations.
-Appendix J Additional Ablations
-J.1 Loss Variant: InfoNCE vs. Margin Hinge
+Appendix 10 Additional Ablations
+10.1 Loss Variant: InfoNCE vs. Margin Hinge
 Replacing our asymmetric InfoNCE (§3.2) with an margin-based hinge loss (margin = 0.2) yields identical maskAP (32.2%) and +0.6 AP50, at the cost of $1.6$× longer training.
@@
 Margin hinge (m = 0.1) 32.2 57.1
-J.2 Ablation: Debias Score Formulation
+10.2 Ablation: Debias Score Formulation
 […]
@@
 Square-root 31.9 56.2
-J.3 Ablation: Negative Sampling Scope
+10.3 Ablation: Negative Sampling Scope
 We evaluate two negative sampling scopes: (i) sampling only within the current mini batch vs. (ii) sampling from a small memory bank of past pixel embeddings (size 10k).

### v4

@@
 Abstract
-Instance segmentation demands costly per-pixel annotations and large models.
-We introduce CAST, a semi-supervised knowledge distillation (SSKD) framework that compresses pretrained vision foundation models (VFM) into compact experts using limited labeled and abundant unlabeled data.
-CAST unfolds in three stages: (1) domain adaptation of the VFM teacher(s) via self-training with contrastive pixel calibration, (2) distillation into a compact student via a unified multi-objective loss that couples standard supervision and pseudo-labels with our instance-aware pixel-wise contrastive term, and (3) fine-tuning on labeled data to remove residual pseudo-label bias.
-Central to CAST is an instance-aware pixel-wise contrastive loss that fuses mask and class scores to mine informative negatives and enforce clear inter-instance margins.
+Instance segmentation demands costly per-pixel annotations and computationally expensive models.
+We introduce CAST, a semi-supervised knowledge distillation (SSKD) framework that compresses pre-trained vision foundation models (VFM) into compact experts using limited labeled and abundant unlabeled data.
+CAST unfolds in three stages: (1) domain adaptation of the VFM(s) via self-training with contrastive calibration, (2) knowledge transfer through a unified multi-objective loss, and (3) student refinement to mitigate residual pseudo-label bias.
+Central to CAST is an instance-aware pixel-wise contrastive loss that fuses mask and class scores to extract informative negatives and enforce clear inter-instance margins.
 By maintaining this contrastive signal across both adaptation and distillation, we align teacher and student embeddings and fully leverage unlabeled images.
-On Cityscapes and ADE20K, our 11$\approx$ 11$\times$ 11 × smaller student surpasses its adapted VFM teacher(s) by +3.4 AP (33.9 vs.
-30.5) and +1.5 AP (16.7 vs.
-15.2) and outperforms state-of-the-art semi-supervised approaches.
+On Cityscapes and ADE20K, our $\approx 11\times$ smaller student improves over its zero-shot VFM teacher(s) by +8.5 and +7.1 AP, surpasses adapted teacher(s) by +3.4 and +1.5 AP, and further outperforms state-of-the-art SSKD methods on both benchmarks.
 Appendix 1 Introduction
 Pixel-level instance segmentation is notoriously expensive: annotating detailed masks can take hours per image, and training state-of-the-art detectors often requires hundreds of GPU hours, putting many applications out of reach [12, 15].
-At the same time, recent advancements in VFMs [22, 20, 43, 19] have significantly elevated the capabilities of computer vision systems, enabling exceptional performance across diverse perception benchmarks [2].
+Recent advancements in vision foundation models (VFMs) [22, 20, 45, 19] have substantially expanded the capabilities of computer vision systems, achieving strong performance across diverse perception benchmarks [2].
 Motivation.
-Despite remarkable achievements, foundation models still cannot serve specific downstream tasks sufficiently well due to two major issues: (1) the heavy computational overhead during deployment making these models impractical for environments with limited resources [37]; and (2) their inherently generic nature, which leads to suboptimal performance on tasks that demand domain specific expertise.
+Despite remarkable achievements, foundation models still cannot serve specific downstream tasks sufficiently well due to two major issues: (1) the heavy computational overhead during deployment making these models impractical for environments with limited resources [38]; and (2) their inherently generic nature, which leads to suboptimal performance on tasks that demand domain specific expertise [28].
 The latter stems from foundation models being optimized to perform well across a wide variety of tasks, rather than being finely tuned for the nuanced requirements of specialized applications [5].
-This challenge is particularly prominent in applications involving outdoor environments, such as autonomous driving, and indoor settings, such as robotic perception [13].
-SSKD for instance segmentation seeks to compress large models into efficient student models by leveraging both limited labeled data and abundant unlabeled images.
-Current distillation methods either treat VFMs as fixed feature extractors with simple pseudo-labeling or focus on coarse semantic tasks, failing to exploit the rich structure of large unlabeled datasets to refine per-pixel predictions.
+This challenge is prominent in applications that involve outdoor environments, such as autonomous driving, and indoor settings, such as robotic perception [13].
+Semi-supervised knowledge distillation (SSKD) for instance segmentation seeks to compress large models into efficient student models by leveraging both limited labeled data and abundant unlabeled images.
+Current distillation methods either treat VFMs as fixed feature extractors with simple pseudo-labeling or focus on coarse semantic tasks, failing to exploit the rich structure of unlabeled datasets to refine per-pixel predictions.
 Consequently, adjacent instances remain poorly separated and accuracy degrades sharply under scarce labels.
-We address these issues by adapting VFMs via self-training to enhance pseudo-label fidelity, and by injecting an instance-aware pixel-wise contrastive loss that leverages unlabeled data to enforce clear inter-instance margins, yielding sharper masks and superior performance in the low-label regime.
+We address these issues by adapting VFMs via self‐training to enhance pseudo-label fidelity, and by injecting an instance-aware pixel‐wise contrastive loss that leverages unlabeled data to enforce clear inter‐instance margins, yielding sharper masks and superior performance in the low‐label regime.
 Status quo.
-Knowledge distillation has evolved from task-agnostic compression [16, 8] to adapting VFMs for downstream tasks.
+Knowledge distillation has evolved from task‐agnostic compression [16, 8] to adapting VFMs for downstream tasks.
 For classification and semantic segmentation, Vemulapalli et al.
 [31] distill a VFM matching its output on an unlabeled transfer set, and SAM-CLIP [32] fuses CLIP and SAM.
-However, neither method targets per-pixel instance masks nor exploits dense self-supervision from the unlabeled pool.
-Pure semi-supervised instance segmentation methods, such as [17, 3] train teachers from scratch, doubling GPU cost, and still produce blurred boundaries under scarce labels.
-To our knowledge, no prior work unifies VFM adaptation, unlabeled data-driven pixel-wise refinement, and extreme student compression for high fidelity instance segmentation.
+However, neither method targets per-pixel instance masks nor exploits dense self‐supervision from the unlabeled pool.
+Pure semi‐supervised instance segmentation methods, such as [17, 3] train teachers from scratch, doubling GPU cost, and still produce noisy masks under scarce labels.
+To our knowledge, no prior work unifies VFM adaptation, unlabeled data‐driven pixel‐wise refinement, and extreme student compression for instance segmentation.
 Contributions. We summarize our main contributions as follows:
-  1. 1.
-    We introduce an instance-aware pixel-wise contrastive loss that fuses mask and class predictions to drive stronger inter-instance separation, and show how to sample negatives efficiently in a instance centric setting.
-  2. 2.
-    We propose CAST, a SSKD pipeline with three phases: (i) self-training a foundation teacher with contrastive calibration, (ii) distilling into a compact student using supervised, pseudo label, and pixel-wise contrastive objectives, and (iii) final supervised finetuning to remove bias thereby unifying supervised, semi-supervised, and self-supervised signals.
-  3. 3.
-    We conduct extensive experiments on Cityscapes and ADE20K, demonstrating that our 11$\approx$ 11$\times$ 11 × smaller student surpass self-trained teacher by +3.4 AP (33.9 vs.
-30.5) and +2.1 AP (16.7 vs.
-14.6), and outperforms state-of-the-art semi-supervised instance segmentation methods under the same data splits, with lower training cost.
+  • We introduce an instance-aware pixel-wise contrastive loss that fuses mask and class predictions to drive stronger inter-instance separation, and show how to sample negatives efficiently in an instance centric setting.
+  • We propose CAST, a SSKD pipeline with three phases: (i) adapting the foundation teacher via self-training with contrastive calibration, (ii) distilling into a compact student using a unified objective that combines supervised, pseudo-label, and pixel-wise contrastive losses, and (iii) supervised fine-tuning to reduce residual bias, unifying supervised, semi-supervised, and self-supervised signals.
+  • We conduct extensive experiments on Cityscapes and ADE20K, demonstrating that our $\approx 11\times$ smaller student improves over its zero-shot VFM teacher(s) by +8.5 and +7.1 AP, surpasses adapted teacher(s) by +3.4 and +1.5 AP, and further outperforms state-of-the-art semi-supervised instance segmentation methods under the same data splits, with lower training cost.
 Appendix 2 Related Work
 Vision Foundation Models.
-Vision foundation models (VFMs)  [22, 20, 25, 41, 4] have revolutionized computer vision through large scale pretraining.
+VFMs [22, 20, 25, 42, 4] have revolutionized computer vision through large scale pre-training.
 In parallel, recent trends focus on combining VFMs to extend their capabilities [26, 45].
 While these models excel in open-set recognition and transfer learning, their computational demands yet hinder edge deployment.
-Recent efforts merge VFMs via distillation: Wang et al [31] unify SAM and CLIP via multi-task learning, while Zhang etal [44] distill CLIP and DINOv2 into a compact model with moderate scale data distillation.
+Recent efforts merge VFMs via distillation: Wang et al.
+[32] unify SAM and CLIP via multi-task learning, while Zhang et al.
+[46] distill CLIP and DINOv2 into a compact model with data distillation.
 We extend these paradigms by leveraging VFMs for instance segmentation, focusing on balancing robustness with computational efficiency.
@@
 Early methods distilled softened logits or intermediate features [16] in a task-agnostic way, while later feature-based approaches capture structured spatial cues (e.g., pixel-wise similarity, channel distributions) [23, 27].
-Modern methods tackle VFMs’ scale and complexity: [28, 38] distills VFMs to impart zero-shot and multimodal capabilities, further multi teacher approaches [18, 39] combine complementary expertise.
+Modern methods tackle VFMs’ scale and complexity: [29, 39] distills VFMs to impart zero-shot and multimodal capabilities, further multi-teacher approaches [18, 40] combine complementary expertise.
 Vemulapalli et al.
 [31] adapt a VFM to the target task and then distill on a large unlabeled set for classification and semantic segmentation.
-Overall with these advances in vision KD We build our method on the insight that a strong teacher (or ensemble of teachers) can guide a lightweight instance segmentation model to high performance.
-our approach explicitly integrates semi-supervised learning and pixel-level contrastive losses for instance segmentation, to focus on bridging the gap between rich representation of VFMs and compact, efficient student networks.
+Building on these advances in vision knowledge distillation, we posit that a strong teacher (or ensemble of teachers) can effectively guide a lightweight instance segmentation model to high performance.
+Our approach explicitly integrates semi-supervised learning and pixel-level contrastive signals for instance segmentation, to focus on bridging the gap between rich representation of VFMs and compact, efficient student networks.
 Semi-Supervised Learning.
-Self-training (or pseudo-labeling) has become a foundational paradigm in semi-supervised learning (SSL), where a model leverages its own predictions with high confidence and iteratively refine itself [35].
+Self‐training (or pseudo‐labeling) has become a foundational paradigm in semi‐supervised learning (SSL), where a model leverages its own predictions with high confidence and iteratively refines itself [36].
 This approach has proven effective across vision tasks, improving image classification performance [36] and boosting object detection accuracy when annotation budgets are tight [21].
-To counteract error accumulation from noisy pseudo-labels [29] use exponential moving average of label predictions, or [6] employ curriculum labeling schemes that gradually incorporate harder examples.
-More recent work applies pseudo-labeling for large pretrained models through targeted finetuning and adaptive pseudo selection strategies [14].
+To counteract error accumulation from noisy pseudo‐labels [30] use exponential moving average of label predictions, or [6] employ curriculum labeling schemes that gradually incorporate harder examples.
+More recent work applies pseudo-labeling for large pre-trained models through targeted finetuning and adaptive pseudo selection strategies [14].
 While many SSL methods focus on classification or detection, several have extended this method to dense prediction tasks [10, 41].
-We study self-training with self-supervised contrastive learning and task-specific adaptation.
-Global contrastive frameworks such as SimCLR [7], MoCo [9], and their detection extensions [34] established the value of large scale visual discrimination learning.
-Further per pixel contrastive approaches [32, 36, 46, 33, 1] have shown promise in retaining spatial sensitivity though they yet conflate pixels from different instances of the same class.
-We extend these advances by synergizing self-training and self-supervised contrastive learning, and introduce a novel instance aware negative sampling strategy designed specifically for the demands of instance segmentation.
+We study self‐training with self‐supervised contrastive learning and task‐specific adaptation.
+Global contrastive frameworks such as SimCLR [7], MoCo [9], and their detection extensions [35] established the value of large-scale visual discrimination learning.
+Further per-pixel contrastive approaches [33, 37, 48, 34, 1] have shown promise in retaining spatial sensitivity though they yet conflate pixels from different instances of the same class.
+We extend these advances by synergizing self-training and self-supervised contrastive learning, and introduce a novel instance-aware negative sampling strategy designed specifically for the demands of instance segmentation.
 Figure 1: CAST framework overview.
-Top: Three-stage pipeline: (1) adapt a pretrained VFM teacher to the target domain via self-training with contrastive pixel calibration; (2) transfer knowledge to a compact student using instance-aware pixel-wise contrastive sampling; (3) fine-tune the student on labeled data to correct residual pseudo-label bias.
-Bottom: Detailed view of stage (2): fused mask and class score maps produce anchor pixels, which are sampled across weak/strong views to form positive and negative pairs; an MLP projects features and computes the pixel-wise contrastive loss.
-Dashed arrows denote no gradient flow; red components are trainable, blue are frozen.
+Top: Three-stage pipeline: (1) adapt a pre-trained VFM teacher to the target domain via self-training with pixel-level contrastive calibration; (2) distill knowledge into a compact student using instance-aware contrastive sampling; (3) fine-tune the student on labeled data to correct residual pseudo-label bias.
+Bottom: Detailed view of stage (2): fused mask and class score maps produce anchor pixels, sampled across weak/strong views to form positive/negative pairs; an MLP projects features for the contrastive loss.
+Dashed arrows denote no gradient flow; red modules are trainable, blue are frozen.
 Appendix 3 Method
 3.1 Overview
-In semi-supervised settings, we are given a small labeled set and a substantially larger unlabeled pool:
+In semi‐supervised settings, we are given a small labeled set and a substantially larger unlabeled pool:
 $\mathcal{D}^{l}=\bigl\{(x_{i}^{l},y_{i}^{l})\bigr\}_{i=1}^{N_{l}}\quad\text{and}\quad\mathcal{D}^{u}=\bigl\{x_{i}^{u}\bigr\}_{i=1}^{N_{u}},\quad N_{u}\gg N_{l},$
@@
 3.3 then details the three stages of the CAST pipeline.
-3.2 Pixel-wise Contrastive Loss
+3.2 Pixel‐wise Contrastive Loss
 Standard supervised and pseudo-label losses enforce correct mask predictions, ignoring pixel-level feature relationships which underutilize unlabeled data and amplify pseudo label noise.
 We therefore inject a self-supervised pixel-wise contrastive loss as an additional supervisory signal on both labeled and unlabeled images, sharpening feature discrimination and regularizing against noisy labels.
-Let $z^{\rm weak},z^{\rm strong}\in\ R^{B\times N\times D}$ be $\ell_{2}$-normalized embeddings from two views of each image, where $B$ is the number of images in one mini batch, $N=h\times w$ the number of pixels, and $D$ the embedding dimension.
+Let $z^{\rm weak},z^{\rm strong}\in\ R^{B\times N\times D}$ be $\ell_{2}$‐normalized embeddings from two views of each image, where $B$ is the number of images in one mini batch, $N=h\times w$ the number of pixels, and $D$ the embedding dimension.
 For each pixel $p\in{1,2,...,N}$ and image index $b\in{1,...,B}$, the corresponding embedding vector is denoted as $z_{b,p}\in\mathbb{R}^{D}$.
@@
 $s^{-}_{b,p,r}=\langle z^{\rm weak}_{b,p},\,z^{\rm strong}_{b^{\prime},q_{r}}\rangle/T,\quad r=1,\dots,R.$
-The pixel-wise contrastive loss is then the standard NT-Xent over all anchors:
+The pixel‐wise contrastive loss is then the standard NT‐Xent over all anchors:
 $\mathcal{L}_{\rm pxl}=-\frac{1}{B\,N}\sum_{b=1}^{B}\sum_{p=1}^{N}\log\frac{\exp(s^{+}_{b,p})}{\exp(s^{+}_{b,p})+\sum_{r=1}^{R}\exp(s^{-}_{b,p,r})}.$
-Debiased Pixel-Level Negative Sampling.
-To mine true inter-instance pairs without quadratic cost, we derive a per pixel sampling distribution by fusing mask and class probabilities.
+Debiased Pixel‐Level Negative Sampling.
+To mine true inter‐instance pairs without quadratic cost, we derive a per pixel sampling distribution by fusing mask and class probabilities.
 Let $M\in\mathbb{R}^{B\times K\times H\times W}$, and $L\in\mathbb{R}^{B\times K\times(C+1)},$ be the model’s mask and class logits respectively.
@@
 $F_{c}[b,p,c]=\sum_{k=1}^{K}P_{m}[b,k,p]\;P_{c}[b,k,c],\quad y[b,p]=\begin{bmatrix}P_{m}[b,1:K,p]\\[6.0pt] F_{c}[b,p,1:C+1]\end{bmatrix}\;\in\;\ R^{\,K+(C+1)}.$
-We score any two pixels $(b,p)\neq(b^{\prime},q)$ by $\tilde{y}$ being $\ell_{2}$-normalized vector of pseudo probability map.
-[…]
+We score any two pixels $(b,p)\neq(b^{\prime},q)$ by $\tilde{y}$ being $\ell_{2}$‐normalized vector of pseudo probability map.
+$s^{\mathrm{deb}}\bigl((b,p),(b^{\prime},q)\bigr)=\max\bigl(0,\;1-\langle\tilde{y}[b,p],\,\tilde{y}[b^{\prime},q]\rangle\bigr),$
 We draw $R$ negatives $\{q_{r}\}$ for each anchor $(b,p)$ by sampling proportional to $s^{\rm deb}$, and then plug these into the NT-Xent denominator of $\mathcal{L}_{\rm pxl}$.
 Theoretical Insight.
-To give a formal rationale for augmenting our pixel-wise contrastive loss, we show that even under a mild negative sampling guarantee, each gradient step on our contrastive term provably increases the expected inter-instance margin.
+To give a formal rationale for augmenting our pixel‐wise contrastive loss, we show that even under a mild negative sampling guarantee, each gradient step on our contrastive term provably increases the expected inter-instance margin.
 Assumption 3.1 (Negative Sampling Guarantee).
@@
 where $\mathcal{D}^{u}=\varnothing$ makes the middle term zero.
-Teacher adaptation. Starting from pretrained weights $\theta_{T}^{0}$, we first fine-tune on the labeled set $\mathcal{D}^{l}$:
-[…]
-We then generate pseudo-labels $\hat{y}_{j}^{u}=f_{\theta_{T}^{\prime}}(x_{j}^{u})$, reset to $\theta_{T}^{0}$and fine-tune on $\mathcal{D}^{l}\cup\{\!(x_{j}^{u},\hat{y}_{j}^{u})\!\}$:
-[…]
-This two-step contrastive calibration yields a specialized teacher whose pseudo-labels are both accurate and spatially consistent for the target domain.
+Teacher adaptation. Starting from pretrained weights $\theta_{T}^{0}$, we first fine‐tune on the labeled set $\mathcal{D}^{l}$:
+$\theta_{T}^{\prime}=\arg\min_{\theta}\;\mathcal{J}\bigl(\theta;\,\mathcal{D}^{l},\varnothing;\,0,\lambda_{\rm pxl}\bigr).$
+We then generate pseudo‐labels $\hat{y}_{j}^{u}=f_{\theta_{T}^{\prime}}(x_{j}^{u})$, reset to $\theta_{T}^{0}$and fine-tune on $\mathcal{D}^{l}\cup\{\!(x_{j}^{u},\hat{y}_{j}^{u})\!\}$:
+$\theta_{T}^{\prime\prime}=\arg\min_{\theta}\;\mathcal{J}\bigl(\theta;\,\mathcal{D}^{l},\mathcal{D}^{u};\,1,\lambda_{\rm pxl}\bigr).$
+This two‐step contrastive calibration yields a specialized teacher whose pseudo-labels are both accurate and spatially consistent for the target domain.
 Knowledge transfer. With calibrated teacher $\theta_{T}^{\prime\prime}$ frozen, student $\theta_{s}$ is trained via the unified objective:
@@
 As a final step, we fine-tune the student on labeled data alone:
-[…]
-This pass removes pseudo-label drift and sharpens decision boundaries for in-domain data.
+$\theta_{s}^{\dagger}=\arg\min_{\theta_{s}^{*}}\;\mathcal{J}\bigl(\theta_{s}^{*};\,\mathcal{D}^{l},\varnothing;\,0,0\bigr),$
+This pass removes pseudo-label drift and sharpens decision boundaries for in‐domain data.
 Appendix 4 Experiments
@@
 ADE20K [49] comprises 20,210 training and 2,000 validation images spanning diverse indoor and outdoor environments, annotated with 150 semantic categories (100 “thing” and 50 “stuff” classes).
-For our semi-supervised experiments, we randomly sample $10\%$ of the Cityscapes training set as labeled and use the remaining images as unlabeled.
-For ADE20K, to reduce computational cost we use stratified sampling to select $20\%$ of the training pool while preserving the original per category distribution, yielding 3,537 images.
-From this subset, 1,000 images (10 per instance class) serve as labeled data and the remaining 2,537 as unlabeled.
-Detailed statistics are provided in Appendix A, and we will release the exact split files alongside our code.
+[…]
 Teacher and Student Architectures.
 Our teacher is a fused ensemble of Grounding-DINO-Large [20] and SAM2-L [25].
-Since the official Grounding-DINO is closed-source, we use its open-source counterpart mm-Grounding-DINO [45].
+Since the SOTA model of Grounding-DINO is closed-source, we use its open-source counterpart mm-Grounding-DINO [47].
 For the student, we pair a DINOv2-S encoder [22] with a DPT-S decoder head [24], followed by a lightweight transformer decoder module in the spirit of Mask2Former [11].
@@
 4, and defer the complete optimizer, learning rate schedules, and other hyperparameters to Appendix B.
-Implementation Details.
-All experiments were conducted on Ubuntu 22.04 with Python 3.10 and PyTorch 2.6.0 (CUDA 12.6).
-Teacher adaptation runs were executed on 2× NVIDIA A100 GPUs, while student training and fine-tuning runs used 2× NVIDIA GeForce RTX 4090 GPUs.
-For instance, a single fine-tuning run of the teacher (Grounding-DINO) on the supervised Cityscapes split required $\approx 3.5$ GPU hours; a single student training run for this dataset took $\approx 17$ GPU hours without the pixel-wise contrastive branch and $\approx 22$ GPU hours with it.
-All reported results in the following tables are averages over three independent runs with different random seeds.
 4.2 Main Results
-We extensively evaluate a range of knowledge distillation (KD) strategies, ranging from purely supervised to state-of-the-art semi-supervised baselines, and benchmark them against our CAST pipeline.
-Table 1 reports maskAP and AP50 on Cityscapes and ADE20K.
-In the teacher adaptation stage (568 M parameters), adding our pixel-level contrastive loss boosts Cityscapes maskAP from 29.8 to 30.5 (+0.7) and AP50 from 54.9 to 56.6 (+1.7); on ADE20K, maskAP rises from 14.6 to 15.2 (+0.6) and AP50 from 23.6 to 24.5 (+0.9).
-These improvements confirm that pixel-wise supervision sharply improves feature discrimination and reduces pseudo-label noise.
-Figure 2 plots key pipeline efficiency metric on a logarithmic scale for both teacher and student models.
-Figure 2: Comparison of Teacher and Student Model efficiency Metrics (Log Scale).
-In the student distillation stage (52 M parameters, just 9 % of the teacher), our semi-supervised model with pixel-level loss achieves 32.2 maskAP and 56.5 AP50 on Cityscapes, outperforming the Guided dist.
-baseline by +1.4 maskAP.
-After fine-tuning, the student reaches 33.9 maskAP (+3.4 over the best teacher) and 58.7 AP50.
-On ADE20K, it records 16.1 maskAP and 27.4 AP50 semi-supervised, and 16.7 maskAP (+1.5) with 28.0 AP50 after fine-tuning, underscoring CAST’s robustness across domains.
-Table 1: Instance segmentation results on Cityscapes and ADE20K. We report teacher-adaptation (568 M params) and student distillation (52 M params) under each method.
-Method Data Regime Cityscapes ADE20K
-maskAP AP50 maskAP AP50
-Teacher Adaptation
-[…]
-Supervised Labeled only $28.7$ $53.4$ $14.2$ $23.5$
-Self-training [35] Labeled + Unlabeled $29.7$ $54.9$ $14.6$ $23.6$
-+ EMA-Augmented [21] Labeled + Unlabeled $29.8$ $54.9$ $14.8$ $23.7$
-+ pixel-level loss (ours) Labeled + Unlabeled 30.5 56.6 15.2 24.5
-Student Distillation
-[…]
-PAIS [17] Labeled + Unlabeled $22.9$ $44.9$ $10.3$ $18.3$
-Guided dist. [3] Labeled + Unlabeled $30.8$ $52.9$ $14.2$ $23.8$
-+ unlabeled KD [30] Unlabeled only $24.4$ $45.6$ $5.1$ $9.3$
-+ labeled+unlabeled KD (ours) Labeled + Unlabeled $30.7$ $54.9$ $14.4$ $25.2$
-+ pixel-level loss (ours) Labeled + Unlabeled 32.2 56.5 16.1 27.4
-Fine-tuned (best) Labeled only 33.9 58.7 16.7 28.0
+We evaluate a range of knowledge distillation (KD) strategies, ranging from purely supervised to state‐of‐the‐art semi‐supervised baselines, and benchmark them against our CAST pipeline.
+Table 1 reports maskAP and maskAP50 on Cityscapes and ADE20K.
+In the teacher adaptation stage (568M parameters), adding our pixel‐level contrastive loss boosts Cityscapes maskAP from 29.7 to 30.5 (+0.8) and maskAP50 from 54.9 to 56.6 (+1.7); on ADE20K, maskAP rises from 14.6 to 15.2 (+0.6) and maskAP50 from 23.6 to 24.5 (+0.9).
+These improvements confirm that pixel‐wise supervision sharply improves feature discrimination and reduces pseudo-label noise.
+In the student distillation stage, our 52M-parameter student (9% of the composite teacher model) achieves 32.2 maskAP and 56.5 maskAP50 on Cityscapes with pixel-level loss, outperforming prior SOTA SSKD models.
+After fine-tuning, the student reaches 33.9 maskAP (+3.4 over the best teacher) and 58.7 maskAP50.
+On ADE20K, it attains 16.1 maskAP and 27.4 maskAP50 in the semi-supervised setting, and improves further to 16.7 maskAP (+1.5) and 28.0 maskAP50 after fine-tuning, underscoring CAST’s robustness across benchmarks.
+Additional ablations under varied label splits are presented in Section 4.
+To compare efficiency, Figure 2 plots key pipeline efficiency metric on a logarithmic scale for both teacher and student models.
+Table 1: Main results on Cityscapes and ADE20K with 10% labeled data. We report teacher adaptation (568M) and student distillation (52M). * denotes adapted methods. Rows in gray are ours.
+Method Data Regime Cityscapes ADE20K maskAP maskAP50 maskAP maskAP50 Teacher Adaptation Zero-shot VFM None (pretrained) $22.0$ $42.3$ $8.1$ $18.2$ Supervised fine-tuning Labeled only $28.7$ $53.4$ $14.2$ $23.5$ Self-training* [36] Labeled+Unlabeled $29.7$ 54.9 $14.6$ $23.6$ Unbiased Teacher* [21] Labeled+Unlabeled 29.8 54.9 14.8 23.7 CAST (teacher adaptation) Labeled+Unlabeled 30.5 56.6 15.2 24.5 Student Distillation Supervised fine-tuning Labeled only $21.1$ $38.7$ $13.9$ $24.2$ PAIS [17] Labeled+Unlabeled $22.9$ $44.9$ $10.3$ $18.3$ Guided dist.
+[3] Labeled+Unlabeled $30.8$ $52.9$ $14.2$ $23.8$ Vemulapalli et al.* [31] Unlabeled only $24.4$ $45.6$ $5.1$ $9.3$ CAST (knowledge transfer) Labeled+Unlabeled 32.2 56.5 16.1 27.4 CAST (student refinement) Labeled only 33.9 58.7 16.7 28.0
+Figure 2: Efficiency comparison (log scale).
 4.3 Empirical Validation
@@
 4.4 Ablation Studies
-We perform a series of ablation experiments on the Cityscapes dataset to isolate the contributions of different components in the CAST pipeline and the pixel-level contrastive loss.
+We perform a series of ablation experiments to isolate the contributions of each component in the CAST pipeline.
+These include analyses of loss functions, training stages, negative sampling strategies, hyperparameters, and student architecture choices.
 Impact of Loss Components.
-In the knowledge distillation stage, the overall loss is a weighted sum of three components: the supervised loss ($\mathcal{L}_{\text{sup}}$), the semi-supervised pseudo-label loss ($\mathcal{L}_{\text{semi}}$), and the pixel-level self-supervised contrastive loss ($\mathcal{L}_{\text{pxl}}$).
-Table 2 presents an ablation study assessing the contribution of each loss term.
-The results indicate that adding the pseudo-label loss substantially improves performance over the supervised baseline, and further inclusion of the self-supervised loss leads to the highest overall performance.
-Table 2: Loss Ablation: Impact of supervised ($\mathcal{L}_{\rm sup}$), pseudo-label ($\mathcal{L}_{\rm semi}$), and pixel-level ($\mathcal{L}_{\rm pxl}$) terms on teacher vs. student maskAP.
-Method $\mathcal{L}_{\rm sup}$ $\mathcal{L}_{\rm semi}$ $\mathcal{L}_{\rm pxl}$ maskAP (%)
-Teacher Student
-(a) Supervised Only 28.7 21.1
-(b) + Pseudo-Label 29.7 30.7
-(c) + pixel-level loss 29.6 27.5
-(d) + (b) + (c) 30.5 32.2
+During distillation, the objective combines three terms: supervised loss ($\mathcal{L}_{\text{sup}}$), semi-supervised pseudo-label loss ($\mathcal{L}_{\text{semi}}$), and pixel-level self-supervised contrastive loss ($\mathcal{L}_{\text{pxl}}$).
+Table 2 shows that adding $\mathcal{L}_{\text{semi}}$ improves student performance from 21.1 to 30.7 maskAP, while further including $\mathcal{L}_{\text{pxl}}$ yields the best result of 32.2 maskAP, confirming complementary benefit.
+Table 2: Ablations on Cityscapes (10% labels). Left: effect of loss terms. Right: effect of CAST stages.
+Method $\mathcal{L}_{\rm sup}$ $\mathcal{L}_{\rm semi}$ $\mathcal{L}_{\rm pxl}$ Teacher Student
+(a) Sup. only 28.7 21.1
+(b) + Pseudo 29.7 30.7
+(c) + Pixel loss 29.6 27.5
+(d) (b)+(c) 30.5 32.2
+(a) Loss ablation
+Variant Teacher Adapt. Distill. Student FT maskAP
+Full CAST 33.9
+No Student FT 32.2
+No Teacher Adapt. 25.7
+Distillation Only 23.8
+No Distill. (Sup.) 21.1
+(b) Stage ablation
+Impact of Training Stages.
+Beyond the contribution of individual loss terms, we further ablate each stage of CAST to justify their necessity.
+Table 2 shows results on Cityscapes (10% labels), where we drop exactly one stage at a time.
+The supervised baseline achieves 21.1 maskAP.
+Adding distillation alone improves this to 23.8 (+2.7), and further adding student fine-tuning raises it to 32.2 (+8.4).
+Without teacher adaptation, performance drops to 25.7, underscoring the need to align the teacher with the target domain.
+The full three-stage CAST pipeline achieves best result of 33.9 maskAP, a +12.8 improvement over baseline.
 Ablation of Negative Sampling via Various Probability Maps.
@@
 The fusion strategy achieves the best results, with 32.2 maskAP and 56.5 AP50.
-(a) Negative Sampling Strategies
-Method maskAP (%) AP50 (%)
+(c) Negative Sampling Strategies
+Method maskAP (%) maskAP50 (%)
 Uniform $29.4$ $50.2$
-Mask-Only $30.6$ $55.0$
-Class-Only $31.1$ $55.3$
+Mask‐Only $30.6$ $55.0$
+Class‐Only $31.1$ $55.3$
 Fusion 32.2 56.5
-(b) Schematic of Sampling Distributions
+(d) Schematic of Sampling Distributions
 Table 3: Ablation of Negative Sampling Strategies on Cityscapes.
-(3) Quantitative results for uniform, mask-only, class-only, and fusion samplers (maskAP and AP50).
-(3) Schematic sketch of the corresponding pixel-level sampling probability distributions.
+(3) Quantitative results for uniform, mask‐only, class‐only, and fusion samplers (maskAP and maskAP50).
+(3) Schematic sketch of the corresponding pixel‐level sampling probability distributions.
 Hyperparameter Sensitivity.
-We evaluate CAST’s sensitivity to three key hyperparameters on Cityscapes: contrastive weight $\lambda_{\rm pxl}$, negatives per anchor $K$, and temperature $T$, by measuring both teacher and student mask AP (%) and AP50 (%).
+We evaluate CAST’s sensitivity to three key hyperparameters on Cityscapes: contrastive weight $\lambda_{\rm pxl}$, negatives per anchor $K$, and temperature $T$, by measuring both teacher and student maskAP (%) and maskAP50 (%).
 Table 4 reports the full sweep.
@@
 For the number of negatives, $K=256$ offers the best trade-off: although $K=512$ yields a slight increase in teacher maskAP (30.9 vs.
-30.5) and AP50 (57.1 vs.
+30.5) and maskAP50 (57.1 vs.
 56.6), and comparable student metrics, the marginal gains saturate relative to the increased sampling cost.
@@
 (a) Encoder Backbone
-Encoder maskAP AP50 Params (M)
+Encoder maskAP maskAP50 Params (M)
 ResNet50 25.5 49.3 24
@@
 (b) Decoder Head
-Decoder maskAP AP50 Params (M)
+Decoder maskAP maskAP50 Params (M)
 FPN 28.9 52.4 18
@@
 Figure 4: Qualitative results on Cityscapes. Guided dist. [3] (top) vs. CAST (bottom).
+Scalability with Labeled Fractions.
+We evaluate CAST under different fractions of labeled data to assess scalability in semi-supervised settings.
+Following the protocol in [3], we train with 5%, 10%, and 30% labeled splits of Cityscapes.
+As shown in Table 6, CAST consistently outperforms prior methods across all fractions.
+At 5% labels, CAST achieves 30.7 AP, far exceeding PAIS (18.0) and Guided Distillation (23.0).
+At 30% labels, CAST reaches 40.4 AP, surpassing the strongest baseline (37.8 from S4M) by +2.6 AP.
+These results demonstrate that CAST remains effective under scarce supervision while scaling gracefully with additional labeled data.
+Table 6: Scalability across label fractions on Cityscapes. Results with 5%, 10%, and 30% labeled data.
+Dataset Fraction Teacher Adapt. Distillation CAST (student) PAIS [17] Guided dist. [3] S4[44]
+5% 29.4 29.2 30.7 18.0 23.0 30.1
+10% 30.5 32.2 33.9 22.9 30.8 33.3
+30% 33.3 38.5 40.4 32.8 35.6 37.8
+Additional ablations, including teacher adaptation variants, loss formulations, sampling scope, and backbone comparisons, are provided in the supplementary material (Section E).
 Appendix 5 Conclusions
-We have introduced CAST, a rigorously designed SSKD pipeline that fuses self-training, instance-aware pixel-wise contrastive learning, and final supervised finetuning to compress large VFMs into compact student experts with comparable performance.
-Empirically, our 11$\approx$ 11$\times$ 11 × smaller student exceeds its adapted teacher by +3.4 maskAP in Cityscapes and +1.5 maskAP in ADE20K, while cutting compute and parameter counts demonstrating that dense contrastive supervision can unlock substantial gains in low-label regimes.
-Our theoretical analysis further guarantees that our negative sampling scheme provably increases inter-instance margins under mild assumptions.
+Figure 5: Performance–complexity radar chart (normalized).
+We have introduced CAST, a rigorously designed SSKD pipeline that fuses self‐training, instance-aware pixel‐wise contrastive learning, and final supervised finetuning to compress large VFMs into compact student experts with comparable performance.
+Empirically, our $\approx 11\times$ smaller student exceeds its adapted teacher by +3.4 maskAP in Cityscapes and +1.5 maskAP in ADE20K, while cutting compute and parameter counts demonstrating that dense contrastive supervision can unlock substantial gains in low‐label regimes.
+Our theoretical analysis further guarantees that our negative sampling scheme provably increases inter‐instance margins under mild assumptions.
 Looking forward, streamlining CAST into a single unified objective, extending its evaluation to diverse domains, and integrating uncertainty quantification will be critical steps toward safe, equitable, and broadly deployable segmentation solutions.
-Figure 5: Performance–complexity radar chart (normalized).
-Broader Impacts.
-This paper focuses on developing high fidelity instance segmentation models under tight annotation and compute budgets, running efficiently on workstation GPUs and making it practical for domains such as autonomous driving, mobile and industrial robotics, augmented reality and agricultural or infrastructure monitoring.
-These safety critical applications are especially vulnerable to biases inherited from generalized VFMs (e.g., underrepresented object classes or ambiguous category definitions, which can lead to degraded performance or unfair outcomes.
-By distilling task-specific expert models, CAST helps reduce these generic failure modes, but careful domain specific validation, bias auditing, and uncertainty quantification remain essential for deployment.
-Limitations.
-While our CAST framework achieves significant performance improvements, the training pipeline comprising multiple stages introduces additional complexity.
-Future work could focus on streamlining these stages and further reducing the training cost.
 Figure 6: Qualitative results on ADE20K.
-Figure 7: Qualitative bias reduction in stage-wise distillation.
-Top row: pseudo-labels generated by the pretrained teacher.
-Bottom row: student predictions after distillation and refinement, demonstrating reduced pseudo-label bias and sharper instance boundaries.
+Figure 7: Qualitative bias reduction in stage‐wise distillation.
+Top row: pseudo‐labels generated by the pretrained teacher.
+Bottom row: student predictions after distillation and refinement, demonstrating reduced pseudo‐label bias and sharper instance boundaries.
+Ethics Statement
+This work does not involve human subjects, private data, or sensitive content.
+All datasets used are publicly available.
+Portions of the manuscript were polished using large language model (LLM) for clarity; this use was limited to text editing and did not affect the research process, experiments, or results.
+Reproducibility Statement
+We provide detailed descriptions of datasets, model architectures, and training procedures to ensure reproducibility.
+All datasets (Cityscapes, ADE20K) are publicly available, and the 5%, 10%, and 30% labeled splits follow established protocols (Section 4).
+Implementation details, including hyperparameters, software environment, and GPU usage, are reported in Section 4.2, with extensive ablations and sensitivity analyses in Appendix E.
+To facilitate further research, we will release our code upon the completion of the anonymous review process.
 Supplementary Material
@@
 Standard validation sets are retained (500 images for Cityscapes, 2 000 for ADE20K).
-Exact image-ID lists will be released with our code.
+Exact image‐ID lists will be released with our code.
 Table 7: Semi-supervised splits used in our experiments.
@@
 Proof Sketch.
-Let $z_{a}$, $z^{+}$ and $\{z^{-}_{r}\}_{r=1}^{R}$ be the unit norm embeddings of an anchor pixel, its positive, and $R$ negatives. Define
+Let $_{a}, z+z^{+}$ and $\{z^{-}_{r}\}_{r=1}^{R}$ be the unit norm embeddings of an anchor pixel, its positive, and $R$ negatives. Define
 $s^{+}=\langle z_{a},\,z^{+}\rangle,\qquad s^{-}_{r}=\langle z_{a},\,z^{-}_{r}\rangle,$
@@
 $\Delta s^{-}$ $=\langle z_{a}^{\prime}-z_{a},\,z^{-}\rangle=\lambda_{\rm pxl}\sum_{r=1}^{R}\alpha_{r}\bigl(\langle z^{+},\,z^{-}\rangle-\langle z^{-}_{r},\,z^{-}\rangle\bigr).$
-By Assumption 3.1, each negative embedding $z^{-}_{r}$ is inter-instance with probability $p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 0$, and intra-instance with probability $1-p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 1$.
+By Assumption 3.1, each negative embedding $z^{-}_{r}$ is inter-instance with probability $p$, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 0$, and intra-instance with probability 1p1-p, in which case $\langle z^{-}_{r},\,z^{+}\rangle\approx 1$.
 Hence
@@
 10.1 Loss Variant: InfoNCE vs. Margin Hinge
-Replacing our asymmetric InfoNCE (§3.2) with an margin-based hinge loss (margin = 0.2) yields identical maskAP (32.2%) and +0.6 AP50, at the cost of $1.6$× longer training.
+Replacing our asymmetric InfoNCE (§3.2) with an margin-based hinge loss (margin = 0.2) yields identical maskAP (32.2%) and +0.6 AP50, at the cost of 1.61.6× longer training.
 This evaluates whether enforcing a fixed positive–negative margin can match or improve upon the performance of InfoNCE.
@@
 10.2 Ablation: Debias Score Formulation
-[…]
-[…]
-[…]
-[…]
-Table 9: Debias Score Formulation Ablation. Original vs. squared vs. square-root debias scores.
+We evaluate three instantiations of the debias score function $s^{\mathit{deb}}$ (§3.2):
+  • Original $s^{\mathit{deb}}$: fusion of mask and class confidences (ours).
+  • $\bigl(s^{\mathit{deb}}\bigr)^{2}$: square each score to amplify high‐confidence negatives.
+  • $\sqrt{s^{\mathit{deb}}}$: take the square root of each score to temper the bias.
+Table 10: Debias Score Formulation Ablation. Original vs. squared vs. square‐root debias scores.
 Score Variant maskAP (%) AP50 (%)
@@
 Squared 32.0 56.3
-Square-root 31.9 56.2
+Square‐root 31.9 56.2
 10.3 Ablation: Negative Sampling Scope
@@
 Scope maskAP (%) AP50 (%)
-Mini-batch only 32.2 56.5
+Mini‐batch only 32.2 56.5
 Memory bank (10k embeddings) 32.7 57.3
-Sampling from a memory bank of 10 k embeddings yields a modest performance gain (+0.5 maskAP, +0.8 AP50) compared to in-batch sampling.
-However, this approach incurs approximately $2.2$× longer training time due to the overhead of maintaining and querying the memory bank.
+Sampling from a memory bank of 10 k embeddings yields a modest performance gain (+0.5 maskAP, +0.8 AP50) compared to in‐batch sampling.
+However, this approach incurs approximately 2.22.2× longer training time due to the overhead of maintaining and querying the memory bank.